# SmartCooper: Vehicular Collaborative Perception with Adaptive Fusion and Judger Mechanism

### 02 Feb 2024 v2

@@ I Introduction
Over the past few decades, autonomous driving perception technology has witnessed remarkable progress, thanks to the rapid advances of computer vision and deep learning techniques. This technology relies predominantly on two types of data sources: image data and point cloud data, both acquired from on-board sensors in real-world scenarios or generated through simulation platforms. In terms of real-world data, researchers have proposed extensive public datasets like KITTI[8] and nuScenes[2]. By using simulation platforms like CARLA[6], simulation datasets like OPV2V[23] and SHIFT[18], usually include a wider range of scenarios and weather conditions, along with comprehensive annotation information, thereby facilitating the studies of autonomous driving perception. Numerous single-vehicle perception methods have been rigorously tested and validated on both real-world and simulation datasets[4],[12],[22]. However, recent developments in communication technology and the advent of CAVs have thrusted collaborative perception into the forefront of autonomous driving research[19]. In contrast to single-vehicle perception, collaborative perception extends the scope of perception and effectively mitigates blind spots[21]. This allows autonomous vehicles to comprehensively perceive their surroundings, leading to more effective decision-making[5]. Specifically, collaborative perception systems facilitate the sharing of perception data among CAVs[14]. Each CAV benefits from a broader sensing range than single-vehicle perception, thus aggregating global and local information to optimize perception results[11]. Existing studies have introduced various collaborative perception methods, including a perception framework based on point cloud features[3] and a collaborative perception framework based on Bird’s-Eye View (BEV), which refers to the top view of scene, and vision[24].
@@
Nonetheless, several challenges persist in current collaborative perception research, particularly when applied to real-world scenarios. Firstly, the data collected by vehicular sensors, comprising image sequences and point cloud data, often pose a significant volume challenge. Regarding image data, contemporary cameras, exemplified by Google’s autonomous vehicles (AVs), are capable of capturing a staggering 750 megabytes of data per second[9]. As for point cloud data, using KITTI[8] as an illustration, even the smallest sequence contains upwards of 10 million data points[13]. The sheer volume of data can easily congest communication channels and potentially result in significant delays or packet loss, thereby restricting effective data sharing among CAVs. Moreover, the real-world communication landscape is far from ideal, and with constrained spectrum resources, , it is difficult to enable real-time data transmissions in Vehicle-to-Vehicle (V2V) networks. Specifically, channel quality fluctuates over time, and the limited bandwidth cannot guarantee consistent real-time and complete data transmission from all vehicles. Some prior studies have already recognized these challenges and initiated efforts to address them. For instance, Hu et al. proposed an efficient communication framework, Where2comm, incorporating an attention-based mechanism to enhance perception performance and mitigate bandwidth pressures[10]. Similarly, Wang et al. introduced a viable approach wherein each agent transmits compressed feature maps, successfully alleviating communication burdens[20]. However, these approaches often assume that channel condition remains time-invariant and fail to account for dynamic channel variations. In contrast, channel-aware methods prioritize the real-time fluctuations in channel conditions among CAVs, adjusting data transmissions accordingly. This dynamic adaptation optimizes communication, ensuring the efficacy of collaborative perception.
@@
Secondly, data compression for CAVs also affects the performance of collaborative perception significantly. However, existing data compression schemes exhibit notable deficiencies. Previous compression methods have typically employed a uniform “fairness scheme” [17], which uniformly applies the same compression ratio to all CAV data. Unfortunately, this fairness scheme overlooks the fact that CAVs in a V2V network contribute perception data differently in data fusion. Some CAVs, situated at some locations, can significantly extend sensing coverage and enhance collaborative perception, while others, with obscured views, provide minimal assistance to data fusion. Moreover, it is crucial to recognize that not all data from CAVs is beneficial. Certain CAVs may be situated at the edge of perception or suffer from severe communication delays, potentially having a detrimental impact on data fusion and consequently reducing the accuracy of collaborative perception. In contrast, a judger mechanism can assess and prioritize CAVs based on their spatial location and sensing coverage. This mechanism can then eliminate low-quality data, enabling efficient CAV data fusion and perception.
@@
Finally, most existing collaborative perception schemes assume that the wireless channel is ideal, ignoring the dynamic characteristics of the channel and the mechanism of filtering the perception data. If the dynamic nature of the channel is taken into account, the compression ratio can be flexibly adjusted and the communication can be optimized. Moreover, the priority can be set adaptively according to the sensing coverage and data quality of CAVs. In contrast to the previous studies, we propose SmartCooper, an adaptive collaborative perception framework with channel-aware communication optimization and judger mechanism to enable efficient CAVs’ data fusion. Our main contributions are summarized as follows.
@@
  • We propose a channel-aware communication optimizing scheme to adapt to the channel state. Specifically, we leverage an adaptive encoder to adjust the compression ratio of data delivered by CAVs according to the real-time channel state, ensuring real-time and high-quality data transmission enabling CAV data fusion.
  • With the OpenCOOD platform[23], we conduct extensive experiments to validate the effectiveness of our proposed scheme by comparing it with the state-of-the-art methods. Under the same channel condition, the communication cost shows a significant reduction by at least 23.10%, and AP@IoU (the average precision of Intersection over Union), which evaluates the accuracy of perception and object detection, is significantly improved by 7.15%.
@@ II-A Collaborative Perception
Due to the limited sensing range of a vehicle’s sensors, CAVs are unable to sense the dangers present in certain hard-to-observe corners. However, collaboration with other CAVs can expand their sensing coverage, thus reducing safety risks. Numerous works on collaborative perception have already been done. Wang et al.[20] proposed V2V collaboration based on point cloud data, while Xu et al.[24] introduced V2V collaboration using 2D image data. Although these studies have demonstrated promising results on test datasets, they often assume an ideal channel environment for V2V collaboration. However, when applying these methods in real-world scenarios, the presence of non-ideal communication channels can pose problems such as high latency and packet loss because of the insufficient network capacity. These challenges can result in suboptimal solutions. Thus, while these approaches have shown potential, the impact of realistic channel conditions needs careful consideration.
@@ II-B Which Vehicle to Communicate
The choice of vehicles to participate in collaboration not only significantly impacts spectrum consumption but also the quality of fused sensing results, thereby affecting the overall performance of cooperation. Therefore, the selection of which vehicles to communicate with is of great importance. Liu et al.[15] proposed a three-stage handshake mechanism for communication selection. Specifically, communicating vehicles are chosen based on matching scores computed between the data from the degraded agent and data from all other normal agents. Hu et al.[10] suggested establishing a spatial confidence map for communicating vehicle selection. However, most of these methods rely heavily on attention mechanisms to build an evaluation system. Assuming a data volume with $N$ communicating vehicles, attention mechanisms implies a time complexity of $O(N^{2})$, a non-negligible cost that could result in additional latency-related issues.
III Proposed Scheme
In this section, we introduce our proposed adaptive collaborative scheme for CAVs. Firstly, in Sec. III-A, we mathematically model the channel conditions in practical communication scenarios, thus defining the adaptive factors required for CAVs in practical settings. Next, in Sec. III-B1, we conduct a concise experiment to illustrate the drawbacks of solely optimizing communication parameters for enhancing collaborative quality. Then, in Sec. III-B2, we introduce a scoring mechanism based on the perceptual gains of the Ego CAV to address the shortcomings mentioned in Sec. III-B1, that is, some CAVs may have detrimental effects on data fusion. Finally, in Sec. III-C we combine the two adaptive strategies and show our comprehensive network architecture.
@@ III-A1 CAV Communication Scheme
In view of the fact that the actual communication situation is not ideal, we propose a channel-aware communication optimizing scheme. We formulate the channel capacity optimization according to CAVs’ actual communication conditions, and propose our adaptive communication scheme. In accordance with the 5G standards by 3GPP[1],[7], CAVs primarily rely on Cellular Vehicle-to-Everything (C-V2X) employing Orthogonal Frequency Division Multiplexing (OFDM). The capacity of each sub-channel $SC_{ij}$ is given by:
$SC_{ij}={{W_{T}}\over N}\log_{2}{{\left(1+{H_{ij}\ {P_{T}}\over{\rho_{n}}}\right)}},$ (1)
where $W_{T}$ denotes the total bandwidth, $N$ the number of sub-channels, $SC_{ij}$ the sub-channel capacity from the $i$th transmitter to the $j$th receiver, $H_{ij}$ the channel gain from the $i$th transmitter to the $j$th receiver, $P_{T}$ the transmit power, and $\rho_{n}$ the noise power spectral density.
@@ III-A2 Channel-aware Adjuster
${\alpha^{*}}_{ij}={{SC_{ij}}\over V_{ij}},$ (2)
where ${\alpha^{*}}_{ij}$ is the compression ratio of the data transmitted from the $i$th transmitter to the $j$th receiver, and $V_{ij}$ is the volume of data transmitted from the $i$th transmitter to the $j$th receiver. Since the volume of data transmitted is often greater than the capacity of a single sub-channel, the compression ratio satisfies: $0\leq{\alpha^{*}}_{ij}\leq 1,{\alpha^{*}}_{ij}\in R^{+}$. Based on this, the adaptive reconstruction structure is designed as follows: we utilize channel-aware adjuster to modulate the compression ratio of the transmitted data for each CAV according to the capacity of the sub-channel, then the encoder transmits the compressed data, and the decoder receives the information transmitted by sub-channel and enables the reconstruction.
@@ III-B Adaptive Judger Mechanism
III-B1 Disadvantage Analysis
The proposed method in Sec. III-A can adaptively allocate communication resources to different CAVs based on the channel environment, thereby enhancing the performance of V2V collaborative perception. However, we have observed that the optimized collaborative data may not always lead to positive gains (i.e., it might potentially degrade the perception performance of the Ego CAV). To illustrate this issue, we conducted a straightforward yet convincing experiment. The results are summarized in Table I.
In our experiment, we employ a test dataset from OPV2V[23] involving a collaborative network of four CAVs. To systematically analyze the impact of participating CAVs on collaborative perception, we increase the number of collaborating CAVs incrementally based on their distance to the Ego CAV. Notably, when the number of CAVs equals 1, it signifies the presence of only the Ego CAV, with no collaborative perception taking place. We conducted tests to measure AP@IoU for the collaborative network under two distinct scenarios: ideal condition and non-ideal condition (Transmissions under non-ideal condition are achieved by adaptive allocating channel resources using the approach outlined in Sec. III-A). Under ideal channel conditions, data transmission does not require any compression. When increasing the number of CAVs, AP@IoU also increases simultaneously. This occurs because different CAVs expand the Ego CAV’s sensing coverage, enhancing its understanding of the surrounding environment. However, under non-ideal channel conditions, an interesting phenomenon emerges when increasing the number of CAVs. Initially, AP@IoU increases, but then it starts to decrease. It indicates that not all CAVs contribute positively in non-ideal conditions. Therefore, we can conclude that optimizing the channel under non-ideal conditions alone is insufficient to achieve optimal collaborative performance.
TABLE I: The collaborative results for different numbers of CAVs under both ideal and non-ideal condition.
@@ III-B2 Scoring & Threshold Scheme
where $s_{cav}$ denotes the sensing coverage of each CAV, and $s_{ego}$ denotes the sensing coverage of ego CAV. The different situations in which the collaborative gain varies with sensing coverage among CAVs are shown in Fig. 2. For the sake of analysis, we assume that the Ego CAV remains stationary. As CAV moves, the variations in the collaborative gain of the Ego CAV exhibit four scenarios: (a) When the distance between the CAV and the Ego CAV is short, the collaborative gain of the Ego CAV increases as the CAV approaches. This trend continues until (b) The intersection of the CAV’s sensing coverage with the Ego CAV’s sensing coverage equals the intersection of the CAV’s sensing coverage with MCSC. In other words, $s_{1}=s_{2}$. (c) As the CAV continues to move, the collaborative gain of the Ego CAV gradually decreases due to most of perception part exceeding MCSC until (d) The CAV moves entirely out of MCSC.
According to Eq. (1) and Eq. (2), the farther a CAV is from the Ego CAV within the MCSC, the fewer channel resources it will be allocated. Additionally, CAVs located far from the Ego CAV obtain low scores based on Eq. (5). These two factors are combined to find CAVs near the vicinity of the MCSC contributing poorer quality and less impactful data to the collaboration with the Ego CAV. Consequently, lack of the consideration of the impact of the detrimental data can lead to interference or even errors. This phenomenon explains the situation observed in cases when an increased number of participating CAVs results in a reduction in collaborative performance in Table I. Hence, guided by experimental data, we have established a threshold based on the experience. This threshold serves to mitigate the adverse impact of detrimental data on collaborative quality. Data falling below the established threshold is discarded outright, while data surpassing the threshold is retained for further processing. It is worth noting that, compared to previous selection strategies based on attention mechanisms, SmartCooper has a lower computational complexity. Assuming there are $N$ CAVs participating in collaboration, the time complexity of the attention mechanism is $O(N^{2})$, whereas SmartCooper has a time complexity of $O(N)$.
(a) Score Increasing
(b) Maximum Score
(c) Score Decreasing
(d) Score equal to zero

### 04 Mar 2024 v3

@@ I Introduction
Over the past few decades, autonomous driving perception technology has witnessed remarkable progress, thanks to the rapid advances of computer vision and deep learning techniques[4][12]. This technology relies predominantly on two types of data sources: image data and point cloud data, both acquired from on-board sensors in real-world scenarios or generated through simulation platforms[16]. In terms of real-world data, researchers have proposed extensive public datasets like KITTI[8] and nuScenes[2]. By using simulation platforms like CARLA[6], simulation datasets like OPV2V[27] and SHIFT[22], usually include a wider range of scenarios and weather conditions, along with comprehensive annotation information, thereby facilitating the studies of autonomous driving perception. Numerous single-vehicle perception methods have been rigorously tested and validated on both real-world and simulation datasets[4],[15],[26]. However, recent developments in communication technology and the advent of CAVs have thrusted collaborative perception into the forefront of autonomous driving research[23][11]. In contrast to single-vehicle perception, collaborative perception extends the scope of perception and effectively mitigates blind spots[25]. This allows autonomous vehicles to comprehensively perceive their surroundings, leading to more effective decision-making[5]. Specifically, collaborative perception systems facilitate the sharing of perception data among CAVs[18]. Each CAV benefits from a broader sensing range than single-vehicle perception, thus aggregating global and local information to optimize perception results[14][10]. Existing studies have introduced various collaborative perception methods, including a perception framework based on point cloud features[3] and a collaborative perception framework based on Bird’s-Eye View (BEV), which refers to the top view of a scene, and vision[28].
@@
Nonetheless, several challenges persist in current collaborative perception research, particularly when applied to real-world scenarios. Firstly, the data collected by vehicular sensors, comprising image sequences and point cloud data, often pose a significant volume challenge. Regarding image data, contemporary cameras, exemplified by Google’s autonomous vehicles (AVs), are capable of capturing a staggering 750 megabytes of data per second[9]. As for point cloud data, using KITTI[8] as an illustration, even the smallest sequence contains upwards of 10 million data points[17]. The sheer volume of data can easily congest communication channels and potentially result in significant delays or packet loss, thereby restricting effective data sharing among CAVs. Moreover, the real-world communication landscape is far from ideal, and with constrained spectrum resources is difficult to enable real-time data transmissions in Vehicle-to-Vehicle (V2V) networks[4]. Specifically, channel quality fluctuates over time, and the limited bandwidth cannot guarantee consistent real-time and complete data transmission from all vehicles. Some prior studies have already recognized these challenges and initiated efforts to address them. For instance, Hu et al. proposed an efficient communication framework, Where2comm, incorporating an attention-based mechanism to enhance perception performance and mitigate bandwidth pressures[13]. Similarly, Wang et al. introduced a viable approach wherein each agent transmits compressed feature maps, successfully alleviating communication burdens[24]. However, these approaches often assume that channel condition remains time-invariant and fail to account for dynamic channel variations. In contrast, channel-aware methods prioritize the real-time fluctuations in channel conditions among CAVs, adjusting data transmissions accordingly. This dynamic adaptation optimizes communication, ensuring the efficacy of collaborative perception.
@@
Secondly, data compression for CAVs also affects the performance of collaborative perception significantly. However, existing data compression schemes exhibit notable deficiencies. Previous compression methods have typically employed a uniform “fairness scheme” [21], which uniformly applies the same compression ratio to all CAV data. Unfortunately, this fairness scheme overlooks the fact that CAVs in a V2V network contribute perception data differently in data fusion. Some CAVs, situated at some locations, can significantly extend sensing coverage and enhance collaborative perception, while others, with obscured views, provide minimal assistance to data fusion. Moreover, it is crucial to recognize that not all data from CAVs is beneficial. Certain CAVs may be situated at the edge of perception or suffer from severe communication delays, potentially having a detrimental impact on data fusion and consequently reducing the accuracy of collaborative perception. Thus, a judger mechanism is needed to assess and prioritize CAVs based on their spatial location and sensing coverage. This mechanism can then eliminate low-quality data, enabling efficient CAV data fusion and perception.
@@
Finally, most existing collaborative perception schemes assume that the wireless channel is ideal, ignoring the dynamic characteristics of the channel and the mechanism of filtering the perception data. If the dynamic nature of the channel is taken into account, the compression ratio can be flexibly adjusted and the communication can be optimized. Moreover, the priority can be set adaptively according to the sensing coverage and data quality of CAVs. In contrast to the previous studies, in this paper, we propose SmartCooper, an adaptive collaborative perception framework with channel-aware communication optimization and judger mechanism to enable efficient CAVs’ data fusion. Our main contributions are summarized as follows.
@@
  • We propose a channel-aware communication optimizing scheme to adapt to the channel state. Specifically, we leverage an adaptive encoder to adjust the compression ratio of data delivered by CAVs according to the real-time channel state, ensuring real-time and high-quality data transmission for CAV data fusion.
  • With the OpenCOOD platform[27], we conduct extensive experiments to validate the effectiveness of our proposed scheme by comparing it with the state-of-the-art methods. Under the same channel condition, the communication cost shows a significant reduction by at least 23.10%, and AP@IoU (the average precision of Intersection over Union), which evaluates the accuracy of perception, is significantly improved by 7.15%.
@@ II-A Collaborative Perception
Due to the limited sensing range of a vehicle’s sensors, CAVs are unable to sense the imminent dangers behind certain hard-to-observe corners. However, collaboration with other CAVs can expand its sensing coverage, thus reducing safety risks. Numerous works on collaborative perception have already been carried out lately. Wang et al.[24] proposed V2V collaboration based on point cloud data, while Xu et al.[28] introduced V2V collaboration using 2D image data. Although these studies have demonstrated promising results on test datasets, they often assume an ideal channel environment for V2V collaboration. When these methods are applied in real-world scenarios, the presence of non-ideal communication channels will cause serious problems such as long latency and high packet loss because of the resulting reduced transmission rate. Thus, while these approaches have shown potential, their effectiveness under realistic channel conditions needs further consideration.
@@ II-B Which Vehicle to Communicate
The choice of vehicles to participate in collaboration not only significantly impacts spectrum consumption but also the quality of fused sensing results, thereby affecting the overall performance of cooperation. Therefore, the selection of which vehicles to participate is of great importance. Liu et al.[19] proposed a three-stage handshake mechanism for communication selection. Specifically, participating vehicles are chosen based on matching scores computed between the data from the degraded agent and data from all other normal agents. Hu et al.[13] suggested establishing a spatial confidence map for participating vehicle selection. However, most of these methods heavily rely on attention mechanisms to build an evaluation system. Assuming a data volume with $N$ participating vehicles, attention mechanisms implies a time complexity of $O(N^{2})$ for communications, a non-negligible cost that could result in additional latency.
III Our Proposed Scheme
In this section, we introduce our proposed adaptive collaborative scheme for CAVs. Firstly, in Sec. III-A, we mathematically model the channel conditions in practical communication scenarios, thus defining the adaptive factors required for CAVs in practical settings. Next, in Sec. III-B1, we conduct a concise experiment to illustrate the drawbacks of solely optimizing communication parameters for enhancing collaborative quality. Then, in Sec. III-B2, we introduce a scoring mechanism based on the perceptual gains of the Ego CAV to address the shortcomings mentioned in Sec. III-B1, that is, some CAVs may have detrimental effects on sensing data fusion. Finally, in Sec. III-C we combine the two adaptive strategies and show our comprehensive network architecture.
@@ III-A1 CAV Communication Scheme
In view of the fact that the actual communication situation is not ideal, we propose a channel-aware communication optimizing scheme. We formulate the channel capacity optimization under CAVs’ actual communication conditions, and propose our adaptive communication scheme. In accordance with the 5G standards by 3GPP[1],[7], CAVs primarily rely on Cellular Vehicle-to-Everything (C-V2X) employing Orthogonal Frequency Division Multiplexing (OFDM). The capacity of each sub-channel ${SC}_{ij}$ is given by:
${SC}_{ij}=\frac{W}{N}\log_{2}{\left[1+\frac{H_{ij}*P_{T}}{\rho_{n}*\left(\frac{W}{N}\right)}\right]},$ (1)
where $W$ denotes the total bandwidth, $N$ the number of sub-channels, ${SC}_{ij}$ the sub-channel capacity from the $i$th transmitter to the $j$th receiver, $H_{ij}$ the channel gain from the $i$th transmitter to the $j$th receiver, $P_{T}$ the transmit power, and $\rho_{n}$ the noise power spectral density.
@@ III-A2 Channel-aware Adjuster
$\alpha^{*}_{ij}=\frac{{SC}_{ij}}{V_{ij}},$ (2)
where ${\alpha^{*}}_{ij}$ is the compression ratio of the data transmitted from the $i$th transmitter to the $j$th receiver, and $V_{ij}$ is the volume of data transmitted from the $i$th transmitter to the $j$th receiver. Since the volume of data transmitted is often greater than the capacity of a single sub-channel, the compression ratio satisfies: $0\leq{\alpha^{*}}_{ij}\leq 1,{\alpha^{*}}_{ij}\in R^{+}$. Based on this, the adaptive reconstruction structure is designed as follows: we utilize channel-aware adjuster to modulate the compression ratio of the transmitted data for each CAV according to the capacity of the sub-channel, then the encoder transmits the compressed data, and the decoder receives the information transmitted over the sub-channel and enables the reconstruction.
@@ III-B Adaptive Judger Mechanism
TABLE I: The collaborative results for different numbers of CAVs under both ideal and non-ideal conditions.
@@
III-B1 Disadvantage Analysis
The proposed method in Sec. III-A can adaptively allocate communication resources to different CAVs based on the channel conditions, thereby enhancing the performance of V2V collaborative perception. However, we have observed that the optimized collaborative data may not always lead to positive gains (i.e., it might potentially degrade the perception performance of the Ego CAV). To illustrate this issue, we conducted a straightforward yet convincing experiment. The results are summarized in Table I.
In our experiment, we employ a test dataset from OPV2V[27] involving a collaborative network of four CAVs. To systematically analyze the impact of participating CAVs on collaborative perception, we increase the number of collaborating CAVs incrementally based on their distance to the Ego CAV. Notably, when the number of CAVs equals 1, it signifies the presence of only the Ego CAV, with no collaborative perception taking place. We conducted tests to measure AP@IoU for the collaborative network under two distinct scenarios: ideal condition and non-ideal condition (Transmissions under non-ideal condition are achieved by adaptively allocating channel resources using the approach outlined in Sec. III-A). Under ideal channel conditions, data transmission does not require any compression. When increasing the number of CAVs, AP@IoU also increases simultaneously. This occurs because different CAVs expand the Ego CAV’s sensing coverage, enhancing its understanding of the surrounding environment. However, under non-ideal channel conditions, an interesting phenomenon emerges when increasing the number of CAVs. Initially, AP@IoU increases, but then it starts to decrease. It indicates that not all CAVs contribute positively under non-ideal conditions. Therefore, we conclude that optimizing the channel under non-ideal conditions alone is insufficient to achieve optimal collaborative performance.
(a) Score Increasing
(b) Maximum Score
(c) Score Decreasing
(d) Score equal to zero
@@ III-B2 Scoring & Threshold Scheme
where $s_{cav}$ denotes the sensing coverage of each CAV, and $s_{ego}$ denotes the sensing coverage of ego CAV. Intuitively, the score captures the sensing area of a collaborative CAV that the ego CAV cannot sense. The different situations in which the collaborative gain varies with sensing coverage among CAVs are shown in Fig. 1. For the sake of analysis, we assume that the Ego CAV remains stationary. As CAV moves, the variations in the collaborative gain of the Ego CAV exhibit four scenarios: (a) When the distance between the CAV and the Ego CAV is short, the collaborative gain of the Ego CAV increases as the CAV approaches. This trend continues until (b) The intersection of the CAV’s sensing coverage with the Ego CAV’s sensing coverage equals the intersection of the CAV’s sensing coverage with MCSC. In other words, $s_{1}=s_{2}$. (c) As the CAV continues to move, the collaborative gain of the Ego CAV gradually decreases due to most of perception part exceeding MCSC until (d) The CAV moves entirely out of MCSC.
According to Eq. (1) and Eq. (2), the farther a CAV is from the Ego CAV within the MCSC, the fewer channel resources it will be allocated. Additionally, CAVs located far from the Ego CAV obtain low scores based on Eq. (5). These two factors are combined to find CAVs near the vicinity of the MCSC contributing poorer quality and less impactful data to the collaboration with the Ego CAV. Consequently, lack of the consideration of the impact of the detrimental data can lead to interference or even errors. This phenomenon explains the situation observed in cases when an increased number of participating CAVs results in a reduction in collaborative performance in Table I. Hence, guided by experimental data, we have established a threshold based on the experience. This threshold serves to mitigate the adverse impact of detrimental data on collaborative quality. Data falling below the established threshold is discarded outright, while data surpassing the threshold is retained for further processing. It is worth noting that, compared to the previous selection strategies based on attention mechanisms, SmartCooper has a lower computational complexity. Assuming there are $N$ CAVs participating in collaboration, the time complexity of the attention mechanism is $O(N^{2})$, whereas SmartCooper has a time complexity of $O(N)$.